# Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity

### v1

@@ Abstract
×$\times\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+\frac{\omega_{0}}{2\left(\Omega-\omega_{0}\right){}^{2}}d_{z}^{2}\bigg{\}}.$ (20)
According to Eq. (17) and Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), it is possible for the excitation rate to surpass the emission rate if the parameters are chosen appropriately.
Discussion.— We have the following comments based on the investigations presented above.
1. It is well known that in a vacuum, an inertial atom in its ground state cannot spontaneously transition to an excited state. However, this is possible for atoms that are uniformly accelerated. From an operational perspective, the spontaneous transition of an accelerated atom from the ground state to an excited state in a vacuum is a direct manifestation of the Unruh effect Davies75 ; Unruh76 ; Fulling73 . The Unruh effect is fundamentally significant but has proven elusive in experimental detection, primarily because the spontaneous excitation is too faint at the accelerations achievable in current experiments. Although it has been shown in Ref. yu24 that the excitation rate of a rotating atom in a vacuum can reach the same order of magnitude as the emission rate when the rotational angular velocity exceeds the atomic transition frequency, resulting in a much higher excitation rate than that of an atom under uniform linear acceleration, the actual excitation rate remains extremely weak, making it difficult to observe in a laboratory setting. To illustrate this explicitly, we take $d=10^{-29}\,\text{Cm}$, $V=10^{-14}\,\text{m}^{3}$, $Q=10^{7}$, $R=50\,\text{nm}$, $\Omega=5\,\text{GHz}$, and $\omega_{0}=10\,\text{MHz}$, satisfying $R\Omega/c\ll 1$. Note that these parameters are experimentally feasible, and consistent with Ref. Lochan20 . With these values, the excitation rate in free space is found to be only $10^{-11}\,{\text{s}}^{-1}$, which is far below the threshold for experimental detection. In stark contrast, if the rotating atom is placed inside a cavity with the normal mode frequency tuned to $\omega_{c}=\Omega-\omega_{0}$, the excitation rate can be amplified to $10^{7}~{}{\text{s}}^{-1}$ according to Eq. (15), which exceeds that in free space by 18 orders of magnitude. Remarkably, this peak value is independent of the atom’s transition frequency and rotational speed, making the choice of parameters less restrictive. The only requirement is the condition $\omega_{c}=\Omega-\omega_{0}$. Meanwhile, the emission rate in this case is negligible, trailing the excitation rate by 9 orders of magnitude. This result highlights the potential for substantial population inversion. Such a dramatic enhancement of the excitation rate could facilitate the experimental detection of the circular Unruh effect and make it a highly feasible goal with current technology.
2. In addition to measuring the cavity-enhanced spontaneous excitation as previously discussed, the circular Unruh effect can also be verified by detecting cavity-enhanced spontaneous emission. By tuning the normal mode frequency of the cavity to $\omega_{c}=\Omega+\omega_{0}$ and using experimentally feasible parameters as before, it is possible to observe emission specifically due to centripetal acceleration, which can reach as high as $10^{7}\,{\text{s}}^{-1}$. More interestingly, by setting the rotational angular velocity and the cavity’s normal mode frequency to $\Omega=2\omega_{c}=2\omega_{0}=5\,\text{GHz}$ while keeping the other parameters unchanged, both the emission and excitation rates can reach as high as $\sim 10^{7}\,\text{s}^{-1}$ simultaneously. This configuration can result in millions of transition events per second for a single centripetally accelerated atom, offering a robust method for experimentally verifying the circular Unruh effect. Such a phenomenon is impossible for an inertial atom, even if its emission rate is higher when $\omega_{c}=\omega_{0}$, because once it emits a photon and transitions to the ground state, it cannot be excited again. In contrast, the continuous transitions between the ground and excited states of a single centripetally accelerated atom serve as compelling evidence for the circular Unruh effect.
3. Based on the discussions above, we argue that manipulating the transition rates and achieving significant population inversion are indeed feasible with current state-of-the-art technologies. Recently, hyperfast rotation of an optically levitated nanoparticle has been achieved by transferring the spin angular momentum of light to the particle’s mechanical angular momentum Reimann2018 ; Li2018 ; Zhang2021 . This progress opens up the promising possibility of attaching an atom to such a rapidly rotating, optically levitated nanoparticle. Notably, the physical parameters used in our numerical investigations ($R=50\,\text{nm}$ and $\Omega=5\,\text{GHz}$) align well with recent experimental developments Zhang2021 , suggesting that the proposed experiments could be practically realizable.
4. Finally, we note that there are recent studies focusing on the transition rates of accelerated atoms within a cavity Scully03 ; Lochan22 ; Lochan20 ; guo24 . Notably, in a recent work Lochan20 where only the emission of excited atoms is investigated, it was demonstrated that the emission rate of a rotating atom due to centripetal acceleration within a cavity can be significantly amplified by the cavity. However, the emission rate reported in Ref. Lochan20 is substantially lower–by 14 orders of magnitude–than the rate found in the current study when using the same parameters. This significant discrepancy arises because, in Ref. Lochan20 , the atoms are assumed to be polarizable only along the direction orthogonal to the plane of rotation. In contrast, our study assumes isotropic polarizability, which is more common in nature. In the non-relativistic limit ($R\Omega/c\ll 1$), the leading term for the emission rate when $\omega_{c}=\Omega+\omega_{0}$ is given by Eq. (11), where the dominant contribution comes from transverse polarization. This rate is approximately $\sim c^{2}/(R\Omega)^{2}~{}(\gg 1)$ times that from axial polarization, making the axial contribution calculated in Ref. Lochan20 negligible.
Summary.— We have investigated the transition rates of a centripetally accelerated atom coupled with fluctuating electromagnetic fields in a vacuum inside a resonant cavity. Our study reveals that the interplay between atomic centripetal acceleration, a form of non-inertial motion, and the confinement of the electromagnetic fields within a high-$Q$ cavity can profoundly alter the radiative properties of atoms. While the maximum cavity-amplified emission rate for inertial atoms cannot be further enhanced by rotation, the spontaneous emission can still be significantly amplified if the atom is far from resonance with the cavity. Notably, the transition rates change dramatically when rotating atoms are placed inside the cavity, enabling significant enhancement of either the emission or excitation rates, individually or even simultaneously.
These findings not only offer a novel approach to controlling the radiative properties of atoms but also pave the way for the experimental verification of the circular Unruh effect using cutting-edge quantum technologies.
Acknowledgements.
This work was supported in part by the NSFC under Grant No. 12075084, No. 12375047, the innovative research group of Hunan Province under Grant No. 2024JJ1006, the Hunan Provincial Natural Science Foundation of China under Grant No. 2023JJ40515, and the Scientific Research Program of Education Department of Hunan Province of China under Grant No. 22B0762.

### v2

@@ II. The excitation rate when $\Omega<\omega_{0}$
$\times\left[\left(1+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}\right)e^{-i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}e^{i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}\right]\;,$ (A22)
$\left\langle\beta\left|B_{l}(t,\mathbf{x})B_{p}\left(t^{\prime},\mathbf{x}^{\prime}\right)\right|\beta\right\rangle=$ $\frac{\hbar}{8\pi\epsilon_{0}V}\int_{0}^{2\pi}d\varphi\int_{0}^{\pi}\sin\theta d\theta\int_{0}^{\infty}d\omega_{k}\rho(\omega_{k})\frac{\omega_{k}}{2c^{2}}\left(\delta_{lp}-\frac{k_{l}k_{p}}{\boldsymbol{k}^{2}}\right)$
$\times\left[\left(1+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}\right)e^{-i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}e^{i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}\right]\;,$ (A23)
$\left\langle\beta\left|E_{l}(t,\mathbf{x})B_{p}\left(t^{\prime},\mathbf{x}^{\prime}\right)\right|\beta\right\rangle=$ $\frac{\hbar}{8\pi\epsilon_{0}V}\int_{0}^{2\pi}d\varphi\int_{0}^{\pi}\sin\theta d\theta\int_{0}^{\infty}d\omega_{k}\rho(\omega_{k})\frac{\omega_{k}}{2c}\epsilon_{lpq}\frac{k_{q}}{\lvert\boldsymbol{k}\rvert}$
$\times\left[\left(1+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}\right)e^{-i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}e^{i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}\right]\;,$ (A24)
[…]
$\times\left[\left(1+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}\right)e^{-i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}+\frac{1}{e^{\frac{\omega_{k}\hbar}{k_{B}T}}-1}e^{i(\omega_{k}t_{-}-\boldsymbol{k}\cdot\boldsymbol{R})}\right]\;,$ (A25)
with $k_{B}$ denoting the Boltzmann constant. By substituting the explicit forms of the field correlation functions into Eq. (2) in the Letter, the leading terms of the transition rates in the non-relativistic limit, i.e., when the linear velocity $v=R\Omega$ is much smaller than the speed of light $c$, can be calculated to be
$\Gamma^{\rm thermal}_{\downarrow}=$ $\frac{\pi}{6V\epsilon_{0}\hbar}\left\{\frac{2\omega_{0}e^{\frac{\omega_{0}\hbar}{k_{B}T}}\rho\left(\omega_{0}\right)}{e^{\frac{\omega_{0}\hbar}{k_{B}T}}-1}d_{z}^{2}+\left[\frac{\left(\omega_{0}+\Omega\right)e^{\frac{\left(\omega_{0}+\Omega\right)\hbar}{k_{B}T}}\rho\left(\omega_{0}+\Omega\right)}{e^{\frac{\left(\omega_{0}+\Omega\right)\hbar}{k_{B}T}}-1}\right.\left.+\frac{\left(\omega_{0}-\Omega\right)e^{\frac{(\omega_{0}-\Omega)\hbar}{k_{B}T}}\rho\left(|\omega_{0}-\Omega|\right)}{e^{\frac{(\omega_{0}-\Omega)\hbar}{k_{B}T}}-1}\right]\left(d_{\rho}^{2}+d_{\phi}^{2}\right)\right\}$
$+\mathcal{O}\left[({R\Omega}/{c}\right)^{2}]\;,$ (A26)
$\Gamma^{\rm thermal}_{\uparrow}=$ $\frac{\pi}{6V\epsilon_{0}\hbar}\left\{\frac{2\omega_{0}\rho\left(\omega_{0}\right)}{e^{\frac{\omega_{0}\hbar}{k_{B}T}}-1}d_{z}^{2}+\left[\frac{\left(\omega_{0}+\Omega\right)\rho\left(\omega_{0}+\Omega\right)}{e^{\frac{\left(\omega_{0}+\Omega\right)\hbar}{k_{B}T}}-1}+\frac{\left(\omega_{0}-\Omega\right)\rho\left(|\omega_{0}-\Omega|\right)}{e^{\frac{(\omega_{0}-\Omega)\hbar}{k_{B}T}}-1}\right]\left(d_{\rho}^{2}+d_{\phi}^{2}\right)\right\}$
$+\mathcal{O}\left[({R\Omega}/{c}\right)^{2}]\;.$ (A27)
where $\mathcal{O}[x^{n}]$ denotes terms on the order of $x^{n}$ or higher that are negligibly small and therefore omitted. By setting $\Omega=0$ in Eqs. (B)–(B), one directly obtains the inertial-case results.
We now present a numerical analysis of the transition rates for a centripetally accelerated atom inside a cavity immersed in a thermal bath at room temperature. These results are obtained using the same parameter set as in Fig. 1 (a), which corresponds to the vacuum case, but with the additional condition of $T=300~\mathrm{K}$. The results are shown in Fig. 3. When the normal mode frequency of the cavity is tuned to $\omega_{c}=\Omega-\omega_{0}$ and $\omega_{c}=\Omega+\omega_{0}$, the emission and excitation rates of the rotating atom reach $10^{11}\;\mathrm{s}^{-1}$, compared with only $10^{-3}\;\mathrm{s}^{-1}$ for an inertial atom, corresponding to an enhancement of $14$ orders of magnitude due to centripetal acceleration. Moreover, this substantial difference between the excitation rates of rotating and inertial atoms in a room-temperature cavity does not rely on fine tuning of the cavity resonance. Even when $\omega_{c}$ is slightly detuned from its optimal value, the excitation rate for the rotating case remains orders of magnitude larger than that for the inertial case. As shown in the insets of Fig. 3 (b), within the range $4.98-5.02\;\mathrm{GHz}$ the excitation rate for rotating atoms remains roughly five orders of magnitude higher, ensuring clear experimental distinguishability. These results demonstrate that rotation can strongly influence the transition rates of atoms inside the cavity even at room temperature, and that this enhancement is experimentally robust.